Suppose that at a busy traffic junction, the probability p of an individual car having an accident is 0.0001. During a certain part of the day, 100 cars pass through the junction. What is the probability of two or more cars being involved in an accident within this period?
1
Expert's answer
2020-06-04T19:10:19-0400

This is a binomial distribution with "n=100,\;p=0.0001."

The binomial distribution can be approximated by the Poisson distribution with "\lambda=np=0.01."

"P(X\ge2)=1-P(X=0)-P(X=1)=e^{-\lambda}(\frac{\lambda^0}{0!}+\frac{\lambda^1}{1!})="

"=1-e^{-0.01}(1+0.01)=0.00004967."